write an equation of the line containing the given point and parallal to the given line.
(8,2);2x-3y=5
The equation of y=?
what does the equation of y equal?
please advise how to solve this type of
question i went with the answer -5 and that was incorrect





Answer by richwmiller(17219)

Write an equation of the line containing the given point and parallel to the given line.
(8,2);2x-3y=5
I will respond to this. Most of what you wrote is unintelligible.
slope of 2x-3y=5
-a/b=-2/-3=2/3
We want a line with the same slope
y=2x/3+b
2=2(8)/3+b
2=16/3+b
2-16/3=b
6/3-16/3=b
-10/3=b
y=2x/3-10/3
or
3y=2x-10